# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v2

@@ III.2.1 Reanalysis of GW170817 alone
We emphasize this simple discussion presents radius constraints directly in terms of conventional nonphysical GW priors on $\Lambda_{i}$, treating each as independent and uniformly distributed. These statistical errors are not informed by any EoS model families, nor even by the requirement that both NS have the same EoS. In addition, the jointly-uniform prior also produces a marginal prior distribution on $\tilde{\Lambda}$ that goes to zero as $\tilde{\Lambda}$ (i.e., $p(\tilde{\Lambda})\propto\tilde{\Lambda}$ at small $\tilde{\Lambda}$). To facilitate visualization and downstream calculation using posterior samples, other studies have reweighted the GW-only posterior distribution for example to be uniform in $\tilde{\Lambda}$ or $R_{1.4}$; see, e.g., Figure 11 of Ref. [2]. These two priors differ substantially: using Eq. (10), a uniform prior in $\tilde{\Lambda}$ (i.e., $P(<\Lambda)=\Lambda/\Lambda_{max}$) corresponds to a cumulative prior distribution that strongly favors the largest radii (i.e., $P(<R_{1.4})\propto R_{1.4}^{6}$). As our calculations employ a directly-interpolated marginal likelihood $L(X)$, not a smoothed posterior, we do not produce such an intermediate data product. Finally, even though the various model-agnostic GW priors used to construct the posterior distributions shown in Figs. 7, 8 and 9 strongly disfavor extremely small NS radii, the GW likelihood itself does not: very small $\Lambda_{i}$ are nearly as consistent with GW170817 observations as $\Lambda_{i}\simeq O(100)$.
III.2.2 Joint EOS inference using GW170817
@@
Gravitational wave measurements, currently favoring smaller radii, are by our prior’s construction more informative than NICER radius constraints. As noted in Figure 1 and the associated discussion, our EoS family and prior enforce a reasonably high radius for low-mass NS leading, through regularity of the mass-radius curve and a need for large maximum mass, to EoS models that favor comparable or larger radii for typical NS. For context, comparing Figures 2 and 1, our prior favors NS radii largely consistent with the radius of J0030, including a wide range between $9~{}{\rm km}$ and $16~{}{\rm km}$ for a fiducial $1.4~{}{\rm M_{\odot}}$ NS. By contrast, as surveyed in Figure 9, GW observations of GW170817 support typically smaller radii for a similar fiducial NS, with little support above $13~{}{\rm km}$ in a direct interpretation of our fiducial GW analyses. As a result and as seen in previous studies, a joint analysis incorporating both galactic pulsars and GW observations finds fiducial pulsar radii comparable to the intersection of these two limits. Figure 10 shows our fiducial calculation incorporating galactic pulsar measurements, GW170817 (with our fiducial prior), and the symmetry energy, following [45]. Using similar parameterizations and observational constraints as a previous study, we find comparable inferences for the nuclear equation of state [45]. Some small differences between our inferences and prior work are expected for technical reasons; for example, we present 90% quantiles in radius at each fixed NS mass, while previous and other works adopt a NS central density prior distribution; see the appendix for further discussion.
Figure 10: Joint EoS constraint from galactic pulsars, symmetry energy, and GW using the $\Gamma$-spectral EoS parameterization. We also provide a benchmark analysis with the same PSRs, symmetry energy, and GW constraint, and uses the same $\Gamma$-spectral parameterization for these constraints [45].
@@
That said, our GW analyses also demonstrate that hidden astrophysical priors have a substantial impact on the inferred NS radius. To characterize this effect, Figure 11 shows a one-dimensional scatterplot of $\ln{\cal E}$ versus $R_{1.4}$ using the three different GW analyses presented in Figure 9. These three analyses draw on identical RIFT marginal likelihood data, differing only in the assumed binary NS population model for mass ratio and spin. We see that the GW signal places an upper bound on the NS radius, and no strong lower bound (base). High mass ratio case (highSpin_lowQ) more strongly and more frequently penalizes higher NS radii, though some configurations with large radii are still highly likely; the overall likelihood shifts to the lowest radii. The comparable masses and positive $\chi_{i,z}$ (pos_highQ) strongly disfavours small radii and is bound on the higher radii showing most of the likelihood in a narrow range of radii on the higher end. This affirms (directly within the context of EoS inference) that astrophysical priors embedded in the GW-only analysis propagate directly into different conclusions about the nuclear EoS.
@@
The pulsar-only likelihoods represented in Figures 2 and 4 and the (multiple) GW-only likelihoods characterized by Figure 11 show that the pulsar and GW measurements may or may not agree, with the amount of tension depending strongly on hidden astrophysical priors. In one scenario, the pulsar/GW results may be in modest tension, consistent with previously presented work. In another scenario, the two may be in good agreement, favoring radii of $13~{}{\rm km}$, if we assume GW170817 is similar to galactic pulsars. Finally, the two approaches could have strong tension, with HESS/GW favoring small radii and NICER favoring large radii, if we assume the kilonova associated with GW170817 favors asymmetric NS binaries. As an example, Figure 12 shows an analysis comparable to Figure 10, excepting only the mass ratio prior ($q<0.6$) applied for GW170817. Despite adopting the same raw GW observation and inference strategy, the same galactic pulsar measurements, and the same symmetry energy, in this scenario we would infer a notably smaller radius for most pulsars, compared to previous work.
@@
III.3 On the maximum neutron star mass and pertinent information that limits it
@@
Despite being quite restricted in form and range, our two headline equation of state families allow for EoS with very large maximum neutron star mass $M_{max}$; see, e.g., Figure 13. The observations of galactic pulsars described above do not identify an upper bound on the neutron star maximum mass, explicitly or implicitly; see, e.g., [55] and references therein. Within our framework, this degeneracy manifests as a broad plateau of the likelihood that covers EoS hyperparameters including such extreme scenarios, as required by this observational constraint; see Figure 4. In our framework these EoS usually exhibit extreme behavior at high density such as their $c_{s}\rightarrow 1$.
Several investigations have proposed theoretically or observationally motivated limits on the EoS which would disfavor these high-$M_{max}$ EoS. In the most extreme approach, Ref. [55] suggests that AT2017gfo’s energetics require a prompt collapse to form a black hole, limiting the neutron star maximum mass to be below $2.16^{+0.17}_{-0.15}~{}{\rm M_{\odot}}$. The bottom panel of Figure 4 shows the specific likelihood we adopt to impose this constraint, following previous work [28].
To illustrate the impact of this suggestion, Figure 14 shows updated results from our fiducial NICER analysis, now incorporating this proposed upper limit. As desired, these constraints disfavor large $M_{max}$. That said, taken with our other constraints, our analysis still allows for the possibility that $M_{\rm max}$ is larger than the median estimate proposed by [55] using this prior information alone, and conversely still disfavors the smallest maximum masses allowed a priori due to the strongly implied minimums on $M_{max}$ through observed NSs.
@@
As a critical side effect, these constraints on $M_{max}$ also inform the EoS over its full dynamic range, shifting and narrowing our posterior distributions in $p(\rho)$ and $c_{s}(\rho)$, particularly at densities well above nuclear saturation density. Further, the soundspeed distribution narrows to higher speeds at high densities in order to compensate for the reduced stiffness at intermediate densities with lower sound speeds and lower pressures to provide with larger pressures that produce stable cores for the substantially compactified star. This results in an M-R that reach Masses of up to $\sim 2~{}{\rm M_{\odot}}$ but does not surpass it substantially.
@@
Figure 14: Impact of maximum mass upper limit: Like Figure 5 – an analysis of galactic pulsar observations and the symmetry energy using the $\Gamma$-spectral EoS – but adding the upper limit on maximum mass proposed in Ref. [55].
III.4 Adding the HESS source
@@
In the main text, we have emphasized EoS inferences using conventional galactic pulsars, explicitly omitting the HESS source J1731 [83] shown in Figure 2. The inferred radius of this source is subject to strong systematic uncertainties [155]. Motivated by previous investigations to assess the impact of this event on EoS inference [156, 157, 158, 159, 160, 161], with Figure 15 we show the results of EoS inference which explicitly include this event, compared with our previous results for galactic pulsars only. Despite the nominal low maximum-likelihood mass and radius deduced for this event, when accounting for the full marginal likelihood we find GW measurements in particular are very consistent with the HESS likelihood. As a result, adding the HESS event produces only incremental changes in the EoS, both as shown here and when adding in GW and other information. In particular, the HESS observation does not seem to require exceptional EoS choices.
@@
Figure 15: Incremental impact of HESS source: Like Figure 5, but adding the HESS source. Except for a modest change at small radius, the overall EoS is only incrementally changed when the HESS source is included, largely at low mass. A benchmark analysis with the same PSRs and symmetry energy and uses the same $\Gamma$-spectral parameterization (also includes GW constraint) is also provided [45].
IV Conclusions

### v3

@@ 3.2.1 Reanalysis of GW170817 alone
We emphasize this simple discussion presents radius constraints directly in terms of conventional nonphysical GW priors on $\Lambda_{i}$, treating each as independent and uniformly distributed. These statistical errors are not informed by any EoS model families, nor even by the requirement that both NS have the same EoS. In addition, the jointly-uniform prior also produces a marginal prior distribution on $\tilde{\Lambda}$ that goes to zero as $\tilde{\Lambda}$ (i.e., $p(\tilde{\Lambda})\propto\tilde{\Lambda}$ at small $\tilde{\Lambda}$). To facilitate visualization and downstream calculation using posterior samples, other studies have reweighted the GW-only posterior distribution for example to be uniform in $\tilde{\Lambda}$ or $R_{1.4}$; see, e.g., Figure 11 of LIGO-GW170817-SourceProperties. These two priors differ substantially: using Eq. (7), a uniform prior in $\tilde{\Lambda}$ (i.e., $P(<\Lambda)=\Lambda/\Lambda_{max}$) corresponds to a cumulative prior distribution that strongly favors the largest radii (i.e., $P(<R_{1.4})\propto R_{1.4}^{6}$). As our calculations employ a directly-interpolated marginal likelihood $L(X)$, not a smoothed posterior, we do not produce such an intermediate data product. Finally, even though the various model-agnostic GW priors used to construct the posterior distributions shown in Figs. 7, 8 and 9 strongly disfavor extremely small NS radii, the GW likelihood itself does not: very small $\Lambda_{i}$ are nearly as consistent with GW170817 observations as $\Lambda_{i}\simeq O(100)$.
3.2.2 Joint EOS inference using GW170817
@@
Gravitational wave measurements, currently favoring smaller radii, are by our prior’s construction more informative than NICER radius constraints. As noted in Figure 2 and the associated discussion, our EoS family and prior enforce a reasonably high radius for low-mass NS leading, through regularity of the mass-radius curve and a need for large maximum mass, to EoS models that favor comparable or larger radii for typical NS. For context, comparing Figures 1 and 2, our prior favors NS radii largely consistent with the radius of J0030, including a wide range between $9\penalty 10000\ {\rm km}$ and $16\penalty 10000\ {\rm km}$ for a fiducial $1.4\penalty 10000\ {\rm M_{\odot}}$ NS. By contrast, as surveyed in Figure 9, GW observations of GW170817 support typically smaller radii for a similar fiducial NS, with little support above $13\penalty 10000\ {\rm km}$ in a direct interpretation of our fiducial GW analyses. As a result and as seen in previous studies, a joint analysis incorporating both galactic pulsars and GW observations finds fiducial pulsar radii comparable to the intersection of these two limits. Figure 10 shows our fiducial calculation incorporating galactic pulsar measurements, GW170817 (with our fiducial prior), and the symmetry energy, following (Miller:2021qha). Using similar parameterizations and observational constraints as a previous study, we find comparable inferences for the nuclear equation of state (Miller:2021qha). Some differences between our inferences and prior work are expected for technical reasons. For example, our inference technique adaptively estimates the posterior distribution, carefully assessing convergence and coverage with an effective sample size. Also, we present 90% quantiles in radius at each fixed NS mass, while previous and other works adopt a NS central density prior distribution; see Section 4 for further discussion.
Figure 10: Joint EoS constraint from galactic pulsars, symmetry energy, and GW using the $\Gamma$-spectral EoS parameterization. We also provide a comparison analysis which adopts the same PSRs, symmetry energy, includes GW170817, and uses the same $\Gamma$-spectral parameterization for these constraints (Miller:2021qha). The calculations differ in their implicit prior along each EOS (our analysis is uniform in mass; the comparison analysis is quadratic in central density, strongly disfavoring low masses and therefore broadening their inherently two-dimensional posterior there); the manner in which GW information is factored in to the analysis (RIFT in our analysis versus a KDE in the comparison, and we omit GW190425); and the inference technique used to sample the space of spectral EOS.
@@
That said, our GW analyses also demonstrate that astrophysical priors have a substantial impact on the inferred NS radius. To characterize this effect, Figure 11 shows a one-dimensional scatterplot of $\ln{\cal E}$ versus $R_{1.4}$ using the three different GW analyses presented in Figure 9. These three analyses draw on identical RIFT marginal likelihood data, differing only in the assumed binary NS population model for mass ratio and spin. We see that the GW signal places an upper bound on the NS radius, and no strong lower bound (base). High mass ratio case (highSpin_lowQ) more strongly and more frequently penalizes higher NS radii, though some configurations with large radii are still highly likely; the overall likelihood shifts to the lowest radii. The comparable masses and positive $\chi_{i,z}$ (pos_highQ) strongly disfavours small radii and is bound on the higher radii showing most of the likelihood in a narrow range of radii on the higher end. This affirms (directly within the context of EoS inference) that astrophysical priors embedded in the GW-only analysis propagate directly into different conclusions about the nuclear EoS.
@@
The pulsar-only likelihoods represented in Figures 1 and 4 and the (multiple) GW-only likelihoods characterized by Figure 11 show that the pulsar and GW measurements may or may not agree, with the amount of tension depending strongly on astrophysical priors. In one scenario, the pulsar/GW results may be in modest tension, consistent with previously presented work. In another scenario, the two may be in good agreement, favoring radii of $13\penalty 10000\ {\rm km}$, if we assume GW170817 is similar to galactic pulsars. Finally, the two approaches could have strong tension, with HESS/GW favoring small radii and NICER favoring large radii, if we assume the kilonova associated with GW170817 favors asymmetric NS binaries. As an example, Figure 12 shows an analysis comparable to Figure 10, excepting only the mass ratio prior ($q<0.6$) applied for GW170817. Despite adopting the same raw GW observation and inference strategy, the same galactic pulsar measurements, and the same symmetry energy, in this scenario we would infer a notably smaller radius for most pulsars, compared to previous work.
@@
3.3 On the maximum neutron star mass and pertinent information that limits it
@@
Despite being quite restricted in form and range, our two headline equation of state families allow for EoS with very large maximum neutron star mass $M_{max}$; see, e.g., Figure 13. The observations of galactic pulsars described above do not identify an upper bound on the neutron star maximum mass, explicitly or implicitly; see, e.g., 2018ApJ...852L..25R and references therein. Within our framework, this degeneracy manifests as a broad plateau of the likelihood that covers EoS hyperparameters including such extreme scenarios, as required by this observational constraint; see Figure 4. In our framework these EoS usually exhibit extreme behavior at high density such as their $c_{s}\rightarrow 1$.
Several investigations have proposed theoretically or observationally motivated limits on the EoS which would disfavor these high-$M_{max}$ EoS. In the most extreme approach, 2018ApJ...852L..25R suggests that AT2017gfo’s energetics require a prompt collapse to form a black hole, limiting the neutron star maximum mass to be below $2.16^{+0.17}_{-0.15}\penalty 10000\ {\rm M_{\odot}}$. The bottom panel of Figure 4 shows the specific likelihood we adopt to impose this constraint, following previous work (Dietrich:2020efo).
To illustrate the impact of this suggestion, Figure 14 shows updated results from our fiducial NICER analysis, now incorporating this proposed upper limit. As desired, these constraints disfavor large $M_{max}$. That said, taken with our other constraints, our analysis still allows for the possibility that $M_{\rm max}$ is larger than the median estimate proposed by 2018ApJ...852L..25R using this prior information alone, and conversely still disfavors the smallest maximum masses allowed a priori due to the strongly implied minimums on $M_{max}$ through observed NSs.
@@
As a critical side effect, these constraints on $M_{max}$ also inform the EoS over its full dynamic range, shifting and narrowing our posterior distributions in $p(\rho)$ and $c_{s}(\rho)$, particularly at densities well above nuclear saturation density. Further, the soundspeed distribution narrows to higher speeds at high densities in order to compensate for the reduced stiffness at intermediate densities with lower sound speeds and lower pressures to provide with larger pressures that produce stable cores for the substantially compactified star. This results in an M-R that reach Masses of up to $\sim 2\penalty 10000\ {\rm M_{\odot}}$ but does not surpass it substantially.
@@
Figure 14: Impact of maximum mass upper limit: Like Figure 5 – an analysis of galactic pulsar observations and the symmetry energy using the $\Gamma$-spectral EoS – but adding the upper limit on maximum mass proposed in 2018ApJ...852L..25R.
3.4 Adding the HESS source
@@
In the main text, we have emphasized EoS inferences using conventional galactic pulsars, explicitly omitting the HESS source J1731 (2022NatAs...6.1444D) shown in Figure 1. The inferred radius of this source is subject to strong systematic uncertainties (2023ApJ...944...36A). Motivated by previous investigations to assess the impact of this event on EoS inference (2023ApJ...958...49S; 2023PhLB..84438062L; 2023PhRvC.108b5806B; 2023A&A...672L..11H; 2023ApJ...949...11J; PhysRevD.109.063017), with Figure 15 we show the results of EoS inference which explicitly include this event, compared with our previous results for galactic pulsars only. Despite the nominal low maximum-likelihood mass and radius deduced for this event, when accounting for the full marginal likelihood we find GW measurements in particular are very consistent with the HESS likelihood. As a result, adding the HESS event produces only incremental changes in the EoS, both as shown here and when adding in GW and other information. In particular, the HESS observation does not seem to require exceptional EoS choices.
@@
Figure 15: Incremental impact of HESS source: Like Figure 5, but adding the HESS source. Except for a modest change at small radius, the overall EoS is only incrementally changed when the HESS source is included, largely at low mass. A comparison analysis with the same PSRs and symmetry energy and uses the same $\Gamma$-spectral parameterization (also includes GW constraint) is also provided (Miller:2021qha).
4 Discussion
Novel results and techniques: Our most novel results systematically demonstrate that the assumed population model for GW170817 will substantially impact the inferred EOS, relative to the effect of an additional pulsar observation – even an extreme observation like HESS J1731. However, in this work we make several additional technical contributions not previously presented elsewhere. First and foremost, we introduced the hyperpipeline technique and provided a distributed and ready-to-use implementation for iterative simulation-based inference with arbitrary user-supplied executables. Next and more specific to EOS inference, building on previous work (2020arXiv200101747W), we systematically demonstrate that astrophysical assumptions about the NS population model are implicitly present whenever using NS observations to constrain the nuclear EOS; how to insert population assumptions self-consistently into calculations of EOS inference; and that the relevant astrophysical population priors generally depend on the formation channel (and observational mechanism) for the observation of interest and not held in common. Third, we perform population inference accounting for both the mass and spin distribution. (Doing otherwise in effect assumes that the mass and spin priors adopted during initial GW parameter inference perfectly reflects the underlying astrophysical population.) Addtionally, in our Appendix, we demonstrate the impact of alternative GW parameter inferences and model systematics. Finally, in a Appendix B, we investigate how alternative mass and distance distributions impact inferences derived from multiple observations, using a novel Fisher-matrix based approach to characterize the population implications of many measurements.
Likelihood model: In this work, we adopt non-constant single-observation likelihoods and marginal likelihoods, adopting consistent probabilistic Bayesian models for the measurement uncertainty associated with each observation. By contrast, as discussed in 2023ApJ...949...11J, many other investigations instead adopt a likelihood which is constant within some allowed domain and zero elsewhere. Their analysis demonstrates that these two approaches can produce surprisingly similar results when applied to the low-dimensional parametric models they employ.
How does our isolated NS model differ from others? For context, other groups have adopted a wide range of different implicit or explicit priors for the neutron star mass distribution. Other priors will inherit different priori mass distributions from their prior on the assumed parameterization $s$. For example, some adopt priors on the neutron star central density, not the neutron star mass (e.g., a quadratic prior (PhysRevD.110.123009; Miller:2021qha) or a log-uniform prior (2019MNRAS.485.5363G) between the maximum central density and some reference density). If instead one follows PhysRevD.110.123009 and chooses $s$ to be the central density and $p(\rho_{c})$ to be some prior on that central density, then the implied mass distribution is proportional to $p(\rho_{c})/dM/d\rho_{c}$. Equivalently, the cumulative distribution function for any individual equation of state satisfies $P(<\rho_{c})=P(<M(\rho_{c}))$, for simplicity assuming a single-branched relationship between mass and central density.
Alternatively, even within the context of nominally uniform mass distributions, others (2020PhRvD.101l3007L) have specifically assumed $p(m|{\cal Y})$ is uniform between a fixed minimum mass and the EoS-dependent maximum TOV mass $M_{max}({\cal Y})$, where $\cal Y$} here and henceforth denotes the EoS or its parameterization. In other words, this model asserts all NSs that can exist under an EoS do exist. In terms of the unit step function $\Theta_{I}(x)=1$ for $x\in I$ (II being the interval) and 0 otherwise, their one- and two-dimensional priors are $p(m)=\Theta_{I}(m)/|I|$ and $p(m_{1},m_{2})=2\Theta_{I}(m_{1})\Theta_{I}(m_{2})\Theta(m_{1}-m_{2})/|I|^{2}$, where $I=[m_{min},M_{max}({\cal Y})]$ is the interval between $m_{min}$ and the EoS-dependent maximum mass. Compared to our fiducial approach, both of these aggressive population models introduce EoS-dependent normalization factors, which weakly penalizes models with large $M_{\rm max}({\cal Y})$ for single observations and strongly penalize them for many observations. Additionally, these aggressive population models should be strongly constrained by the absence of massive neutron stars; see, e.g., Alsing:2017bbc; 2020PhRvD.102f4063C and references therein. Though a joint analysis was not performed in this work, 2020arXiv200101747W demonstrated that, if not known a priori, the neutron star mass distribution and EoS must be jointly inferred, to avoid biasing the deduced EoS.
Implicit mass priors and the EOS posterior: Many other studies present mass-radius distributions nominally for the nuclear EOS alone. These discussions of the NS mass-radius distribution adopt priors along the family of stable NS by, for example, adopting a prior on the NS central density; see, e.g., 2019MNRAS.485.5363G; PhysRevD.110.123009; Miller:2021qha. These analyses are in effect adopting a prior on the NS mass distribution. As a result, in order to replicate their results, we must present our results in a comparable fashion. Specifically, rather than construct a 90% credible interval for the NS radius at each fixed mass, we must instead construct a two-dimensional marginal distribution in both mass and radius. This two-dimensional posterior distribution will explicitly depend on the choice of NS mass distribution (or equivalently NS central density distribution).
Rather than invoke an implicit mass distribution, we perform Bayesian inference on the EOS posterior alone and represent it accordingly. Bayesian inference frameworks provide a likelihood and hence (with a prior) probability for each equation of state function. We can therefore interpret observations as providing information about a credible interval on the space of functions: for example, deducing a volume of EoS which contain 90% of all (joint) posterior probability. Figure 2 shows our 100 percentile prior EoS distribution. This interpretation, however, conflicts with the customary approach to rendering and reporting conclusions about the EoS, which are presented as marginal distributions for (uncorrelated) point measurements: for example, the distribution of pressure at a given density, or the radius of NS of a fiducial mass. As highlighted in previous work (Landry:2020vaw), the superposition of point measurements neglects important information including correlations between different densities (for $p(\rho)$) or masses (for $R(M)$). As a concrete example for any two densities $\rho_{i}$, the posterior pressures $\ln p(\rho_{1}),\ln p(\rho_{2})$ will be correlated by measurements, above and beyond any correlations introduced by our prior. Nonetheless, in this work to simplify comparison with previously-reported results, except for Figure 2 we render our conclusions using this point-measurement approach. Thus for example a confidence interval on $p(\rho)$ follows by taking all weighted EoS realizations $(p_{k}(\rho),w_{k})$ and, for each reference density $\rho_{i}$, constructing a confidence interval for $p(\rho_{i})$ using the weighted sample at that density. Similarly but not identical are confidence intervals on $\rho(p)$, constructed similarly but using the inverse relationship.
Assumptions disguise sources of tension: The individual underlying inferences involve many often subtle steps which disguise their implicit and explicit priors, particularly when combining very heterogeneous sources of external information. Unsurprisingly and closely related, looking more closely at individual measurements and their systematics suggests tension between these different observational approaches to the EoS. For example, several studies have pointed out that the amount and nature of material ejected from a merger like GW170817 is inconsistent with the matter needed to reproduce the details of the kilonova emission (2021ApJ...906...98N; 2020ARNPS..7013120R; 2021ApJ...910..116K; PhysRevResearch.5.013168; Ristic22; 2025arXiv250312320R). Conversely, several investigations have wholeheartedly embraced combining all available measurements to produce overall estimates, despite the incremental impact most of these measurements have on overall inferences, owing to large statistical measurement errors (let alone poorly understood systematics); see, e.g., the discussion in (Dietrich:2020efo). Finally, as highlighted by previous work (2017ApJ...844..156R; 2021PhRvL.126f1101A; 2021ApJ...908..103R; 2022PhRvD.105d3016L), the relative impact of all measurements derives strongly from their explicit prior assumptions, both on the small and large scale. For example, compared to theoretically well-motivated priors (2021ARNPS..71..403D; 2021ARNPS..71..433L), adopting an extremely conservative phenomenological approach with much wider priors enables individual measurements to strongly rule out individual realizations (e.g., PhysRevD.110.123009). Conversely, a too-tightly-parameterized approach could conceivably miss subtle degeneracies associated with putative phase transitions; see, e.g., 2023PhRvL.130t1403R and references therein.
Unexpected features in source parameter inference with mass ratio constraints: Practitioners of GW source parameter inference often worry that posteriors which exhibit upper or lower limits other than that parameters’ natural limit (e.g., $m_{2}/m_{1}\in[0,1]$) may reflect user error. By construction, however, our source parameter inferences for GW170817 conditioned on our two alternative prior distributions (b) and (c) described in Section 2.5 have either upper or lower limits on $m_{2}/m_{1}$ imposed by our adopted prior. In other words, the sharp edges in the posterior distributions seen in Figure 7 as desired reflect the appropriate prior limits adopted in each analysis respectively.
5 Conclusions